# Kinetic energy in random recurrent neural networks

### v1

@@ III.4 Arc Length parametrization of RNN activity
In this section, we study the arc length of RNN trajectories. As time flows, the activity travels along a curve in the manifold embedded in an ambient space of $\mathbb{R}^{N}$. The arc length is thus the cumulative distance traveled from $t_{0}$ to $t$, depending on an integral of speed (related to the kinetic energy). We can compare the results with predictions from DMFT. The arc length $s(t)$ from an initial time $t_{0}$ to time $t$ is thus defined as follows:
@@
As shown in Fig. 6 (a,c), we estimate the arc length of RNNs at various coupling strengths $g=1.2,1.3,1.5$, with network sizes of $N=2000$ or $N=5000$. Each trajectory is averaged over 100 independent realizations of RNN dynamics. The results demonstrate that, in the chaotic regime, the arc length increases with time, and this growth rate becomes larger for stronger coupling strengths. In the long-time limit, the system reaches a steady state, resulting in a rough linear growth of arc length with time.
In Fig. 6(b,d), we perform a linear fit to the arc length trajectory over the final interval $t\in[95,100]$.The resulting slope $d_{s}$ is extracted for different values of $g$. According to the theoretical framework presented earlier, in the steady state, this slope is expected to match the square root of the system’s kinetic energy $\sqrt{\Gamma_{0}}$, predicted by DMFT. We thus compare the fitted slope $k_{s}$ of the simulated arc length with the DMFT prediction $k_{d}=\sqrt{\Gamma_{0}}$, where $\Gamma_{0}$ is computed via Eq. (13) using the self-consistent solution of $\mathcal{F}(\Delta_{0};\Delta_{0})=0$ . As shown in Fig. 6(b,d), the error bars represent the standard deviation over 100 evaluations of Eq. (13), confirming that the arc length is directly related to the kinetic energy of the system. The agreement between the fitted slopes and DMFT predictions further validates this connection. The discrepancies between simulation and theory arise in the case of $g$ close to $g_{c}$, probably due to numerical precision in both solving the self-consistent equation and simulating the RNN dynamics.
Figure 6: Arc length of RNN trajectories and comparison with DMFT results. For simulations, two network sizes are considered. In simulations, $dt=0.1$ being the time unit. (a, c) Arc length computed for RNNs with coupling strengths $g=1.2,1.3,1.5$, averaged over $100$ realizations. (b, d) The slope $k_{s}$ of arc length obtained by a linear fitting over the last $50$ time steps is compared with the DMFT-predicted slope $k_{d}=\sqrt{\Gamma_{0}}$ from Eq. (13). Error bars represent standard deviations across $100$ trajectories. For focusing on the slope, the intercept $c$ in the linear fit $s(t)=kt+c$ is set to zero in the plot.
@@ IV Conclusion
In this work, we study the kinetic energy level of the high-dimensional neural dynamics in the long-time limit. Supposing that the kinetic energy in the stationary limit is an order parameter, a continuous chaos transition is identified, consistent with previous studies with different order parameters [16, 3]. Furthermore, the scaling behavior of the kinetic energy is theoretically analyzed around the critical point, and a power law is identified. By solving the DMFT equations, we also calculate the single-time probability distribution of activity underlying the non-zero kinetic energy, which is in perfect agreement with experiments. One can also study the arc length traveled along the chaotic manifold, which can be further analyzed in the context of computational tasks such as perception or other types of brain states [19, 20]. In future works, it is interesting and important to determine how the dynamics are confined in the phase space to yield a given level of kinetic energy. Our recent work gave a slow-dynamics framework [13], which must be further generalized to a more challenging setting as used in this paper.
@@ Acknowledgements.
This research was supported by the National Natural Science Foundation of China for Grant numbers 12475045 and 12122515, and Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (No. 2022B1212010008), and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023B1515040023).
@@ Appendix A Details of DMFT Derivation
A formal solution to Eq.(2) reads:
$x(t)=\int_{-\infty}^{t}\mathrm{d}se^{-(t-s)}\eta(s).$ (24)
@@
where the function $\Phi(x)=\int_{0}^{x}\mathrm{d}y\phi(y)$ is introduced, and using the Price’s theorem: $\frac{\partial C_{\Phi}(\Delta;\Delta_{0})}{\partial\Delta}=C_{\phi}$ (the subscript indicates the non-linear function used to compute the correlation function, e.g., $\phi=\tanh$ in this paper), we get Eq. (10).
@@ Appendix B Kinetic energy in the limit $g\to 1^{+}$
To have Eq. (38), we repeatedly used Price’s theorem [Price-1958].
@@ Appendix C Algorithmic implementation of the DMFT equation
We add here the details about the implementation of the single-degree-of-freedom DMFT equations, as shown in Algorithm 1. For more implementation details of the dynamical mean-field theory, readers may refer to a pedagogical introduction [2].

### v2

@@ III.4 Arc length parametrization of RNN activity
As illustrated in Fig. 5, representative trajectories of the RNN and the corresponding DMFT dynamics are shown for different values of the coupling strength $g$. The network activity $\bm{x}(t)\in\mathbb{R}^{N}$ evolves on the chaotic attractor with fluctuations increasingly pronounced as $g$ is increased. This is related to the growth of the kinetic energy, which can be linked to the arc length of trajectories studied in this section. As time flows, the activity travels along a curve on the manifold embedded in an ambient space of $\mathbb{R}^{N}$. The arc length is thus the cumulative distance traveled from $t_{0}$ to $t$, depending on an integral of speed (related to the kinetic energy). We can compare the results with predictions from DMFT. The arc length $s(t)$ from an initial time $t_{0}$ to time $t$ is thus defined as follows:
@@
Figure 7: Arc length of RNN trajectories and comparison with DMFT results. For simulations, two network sizes are considered. In simulations, $\mathrm{d}t=0.1$ is the time unit. (a, c) Arc length computed for RNNs with coupling strengths $g=1.1,1.3,1.5$ for two different network sizes, averaged over $100$ realizations. (b, d) The slope $k_{s}$ of arc length obtained by a linear fitting over the last $200$ time steps ($\Delta t=200\times dt$) is compared with the DMFT-predicted slope $k_{d}=\sqrt{\Gamma_{0}}$ from Eq. (13). Shaded regions represent standard deviations across $100$ trajectories. Focusing on the slopes, we fix the intercept $c$ in the linear fit $s(t)=kt+c$ to the height of the circle marker in (a) or (c).
Figure 7 (a,c) shows the arc length obtained from direct RNN simulations for different coupling strengths $g=1.2,1.3,1.5$ and for network sizes $N=2000$ and $N=8000$. Each curve (dashed line) is averaged over $100$ independent realizations. In the chaotic regime, $s(t)$ grows monotonically with time, and its growth rate increases with $g$, reflecting the growth of dynamical fluctuations. After a transient stage, the system reaches a statistically stationary state in which the arc length exhibits an approximately linear growth $s(t)\simeq k_{s}\,t$ (subtracted by a baseline), indicating a constant average speed in state space.
In Fig. 7 (b,d), we perform a linear fit to the arc length trajectory over the final interval $t\in[180,200]$. The resulting slope $d_{s}$ is extracted for different values of $g$. According to the theoretical framework presented earlier, in the steady state, this slope is expected to match the square root of the system’s kinetic energy $\sqrt{\Gamma_{0}}$, predicted by DMFT. We thus compare the fitted slope $k_{s}$ of the simulated arc length with the DMFT prediction $k_{d}=\sqrt{\Gamma_{0}}$, where $\Gamma_{0}$ is computed via Eq. (13) using the self-consistent solution of $\mathcal{F}(\Delta_{0};\Delta_{0})=0$ . As shown in Fig. 7 (b,d), the shaded regions represent the standard deviation over $100$ evaluations of Eq. (13) and RNN simulations, confirming that the arc length is directly related to the kinetic energy of the system and the estimation error is suppressed by increasing network size. The agreement between the fitted slopes and DMFT predictions further validates this connection. The discrepancies between simulation and theory arise in the case of $g$ close to $g_{c}$, probably due to numerical precision in both solving the self-consistent equation and simulating the RNN dynamics.
To further elucidate the origin of slight deviations between simulations and DMFT predictions, we study the temporal fluctuations of the instantaneous kinetic energy $\Gamma_{tt}$ for different network sizes at fixed coupling $g=1.5$, as shown in Fig. 8. While the time-averaged value converges to the DMFT prediction, finite-size networks exhibit strong temporal fluctuations that decrease with increasing $N$. These fluctuations translate into variations of the instantaneous speed and hence into fluctuations in the measured arc-length growth rate (Fig. 7). We expect that in the thermodynamic limit, such fluctuations vanish, and the linear relation between arc length and the square root of kinetic energy becomes exact.
Figure 8: The fluctuations of kinetic energy $\Gamma_{tt}$ as a function of time in the stationary regime for different network size N=2 000,6 000,30  [^0] N=2\,000,6\,000,30\,000. The synaptic gain is set to $g=1.5$. Shaded regions indicate the range between the maximum and minimum values obtained from the selected trajectories for different network sizes. The dash-dotted line indicates the DMFT prediction.
@@ IV Conclusion
In this work, we study the nature of high-dimensional neural dynamics in terms of kinetic energy within the framework of dynamical mean-field theory. We characterize the chaotic regime through the stationary mean kinetic energy, which quantifies the typical speed of network trajectories in phase space in the stationary regime [8]. By solving the DMFT equations in the stationary limit, we show that the average kinetic energy emerges continuously at the well-known transition to chaos and exhibits a cubic growth close to the critical coupling. This scaling behavior is consistent with numerical simulations and provides a coherent description of the onset of chaos, being of the same critical exponents as linear dimensionality of chaotic attractors and topological complexity [25, 19].
By comparing the stationary activity statistics of the original random recurrent network with those generated by a gradient-descent dynamics with the same kinetic energy, we clarify through finite-size network simulations that the activity statistics are approximately the same, but in terms of geometric organization, they are different, highlighting the necessity to derive theoretical arguments in the thermodynamic limit.
The arc length the neural trajectory flows along the chaotic manifold can also be linked to the stationary kinetic energy predicted by DMFT. In particular, the arc length grows linearly in time in the stationary regime, with a rate determined by the stationary kinetic energy. This establishes a direct link between microscopic dynamics and the emergent properties of chaotic attractors in high-dimensional neural systems. From a computational perspective, the arc-length-based characterization of chaotic manifolds may provide useful insights toward understanding neural state representations in cognitive circuits [31, 32, 33].
Our work reveals that the kinetic energy is a key macroscopic observable in understanding the nature of high-dimensional chaos. On one hand, the standard DMFT can determine the steady state behavior of the chaos; on the other hand, the optimization of kinetic energy leads to fixed point statistics, which can be used to clarify the relationship between equilibria and the chaotic attractors, by taking inputs from DMFT. It is thus promising to further explore this kinetic-energy-centric perspective in high-dimensional neural dynamics, linking dynamics geometry, non-equilibrium phase transition, and information processing (see a recent work on neural Langevin machine [34]).
@@ Acknowledgements.
This research was supported by the National Natural Science Foundation of China for Grant numbers 12475045, and Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (No. 2022B1212010008), and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023B1515040023).
@@ Appendix A Details of DMFT Derivation
This part provides a self-contained derivation of known DMFT results for completeness and pedagogical clarity. A formal solution to Eq. (2) reads:
$x(t)=\int_{0}^{t}\mathrm{d}se^{-(t-s)}\eta(s).$ (30)
@@
where the function $\Phi(x)=\int_{0}^{x}\mathrm{d}y\phi(y)$ is introduced, and using the Price’s theorem [29]: $\frac{\partial C_{\Phi}(\Delta;\Delta_{0})}{\partial\Delta}=C_{\phi}$ (the subscript indicates the non-linear function used to compute the correlation function, e.g., $\phi=\tanh$ in this paper), we get Eq. (10).
@@ Appendix B Kinetic energy in the limit $g\to 1^{+}$
To have Eq. (44), we repeatedly used Price’s theorem [29].
@@ Appendix C Algorithmic implementation of the DMFT equation
We add here the details about the implementation of the single-degree-of-freedom DMFT equations, as shown in Algorithm 1. For more implementation details of the dynamical mean-field theory, readers may refer to a pedagogical introduction [5]. More efficient methods are presented in previous works [2, 36].